OpenStudy (anonymous):

(X-3)^2+(y-5)^2=50 center and radius please help!

OpenStudy (whpalmer4):

formula for a circle: \[(x-h)^2 + (y-k)^2 = r^2\]gives you a circle with center at \((h,k)\) and radius \(r\)

OpenStudy (anonymous):

centre (3,5) radius \[\sqrt{50}\] equation is of the (x-a)^2+(y-b)^2=r^2 centre (a,b) radius r